OpenStudy (anonymous):

Find thr sixth term of an arithmetic sequence with tl=2 and tn=tn-1+4

OpenStudy (lgbasallote):

tn = tn-1+4??

OpenStudy (anonymous):

t1 = 2 and tn = tn-1 + 4.

OpenStudy (lgbasallote):

tn = tn-1 + 4?? how is that possible?? are you sure you copied it right?

OpenStudy (anonymous):

Find the sixth term of an arithmetic sequence with t1 = 2 and tn = tn-1 + 4.

OpenStudy (anonymous):

Lol yeah

OpenStudy (diyadiya):

\[\Large t_n=t_{n-1} + 4\] Like this ?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Find S10 for 1 + 4 + 7 +

OpenStudy (anonymous):

Think about it this way: to get to the sixth term from the first term, you're going to need to add the number \(4\) a total of \(6-1=5\) times.